# Kernel vs. Kernel: Exploring How the Data Structure Affects Neural Collapse

### v2

@@ 5.2 Limiting NTK
Under Assumptions 1-3, let $\phi(\cdot)$ be the ReLU activation. Denote by $\mathbf{H}_{GP},\mathbf{H}_{NTK}$ the features associated with NNGP $Q_{GP}^{(1)}$ and NTK $\Theta_{NTK}^{(2)}$, respectively. Then:
@@
Appendix D presents the proof by calculating the expected values of $Q_{GP-ReLU}^{(1)}(\mathbf{x}^{c,i},\mathbf{x}^{c^{\prime},j})$ and employing Theorem 4.1. For a better understanding of the result, consider the balanced class scenario with $n_{1}=n_{2}=N/2$. This gives us: $\mathbb{E}\left[\mathcal{N}\mathcal{C}_{1}(\mathbf{H}_{GP/NTK})\right]=2(\sigma_{1}^{2}+\sigma_{2}^{2})/(\mu_{1}-\mu_{2})^{2}+\Delta_{h.o.t}$, which intuitively captures the sum of the within-class variance of $\mathbf{X}$ in the numerator and the between-class variance in the denominator of the first term.
@@
$\bullet$ Erf Activation. We present a similar analysis with the NNGP and NTK with Erf activation in Appendix E (as the terms involved in the formulation are relatively complex than the ReLU case). In the case of ReLU with balanced data, observe that the numerator (corresponding to $\Sigma_{W}(\mathbf{H}_{GP/NTK})$) solely depends on $\sigma_{c}^{2}$, however, for Erf activation, our analysis shows a dependence on terms $\propto\sigma_{c}^{2}\mu_{c}^{-6}$ i.e, it depends on inverse of higher powers of class means $\mu_{c}$ as well (see (92) in Appendix E). Similar analysis for $\boldsymbol{\Sigma}_{B}(\mathbf{H}_{GP/NTK})$ with Erf in (100) shows a dependence on terms $\propto\sigma_{c}^{2}\mu_{c}^{-4}$. Importantly, under Assumptions 1-3, we show similar values of the expected NC1 metric for NNGP and NTK even for the Erf activation, which are smaller than the ReLU case.
@@
Remark. These results reveal the effect of the activation function on the NC1 metric when $d_{1}\to\infty$. Especially under Assumptions 1-3, the Erf-based kernels reflect a larger extent of ‘variability collapse’ (NC1) of the hidden layer post-activations of our 2L-FCN (both at initialization via NNGP and during training via NTK). Additionally, they indicate that the expected $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$ of the NTK closely approximates the NNGP counterparts on our 1-D Gaussian dataset. Perhaps surprisingly, this shows that NTK does not represent more collapsed features than NNGP, despite being associated with NN gradient-based optimization. Namely, we have established another result that shows that training in the lazy regime provably deviates from the practical feature learning of NNs [41, 42, 43, 44, 45].
5.4 Experiments with High-Dimensional Gaussian Data
Setup: We conduct experiments on datasets with varying sample sizes and input dimensions to verify our theoretical results and show that insights generalize (e.g., beyond $d_{0}=1$). For $C=2$, a dataset size $N$ chosen from $\{128,256,512,1024\}$, and input dimension $d_{0}$ chosen from $\{1,2,8,32,128\}$, we create the data vector and label pairs as follows:
The vectors and labels from the dataset can then be arranged into the matrix form (as described in the setup) for analysis. The sampling procedure is repeated 1010 1010 times for each $(N,d_{0})$ [^2] The code is available at: https://github.com/kvignesh1420/shallow_nc1 .
Observations: Figure 2 illustrates the mean and standard deviation (std) of $\log_{10}(NC1(\mathbf{H}))$ for the post-activation NNGP kernel $Q^{(1)}_{GP}$ and NTK $\Theta^{(2)}_{NTK}$ for Erf and ReLU activations. For the low-dimensional case of $d_{0}=1$ and Erf activation, observe from Figure 2(a) that $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$ has a small value of $\approx 10^{-2.1}$. On the contrary, Figure 2(c) illustrates that for ReLU, $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$ is more than an order of magnitude larger ($\approx 10^{-0.95}$) than the former. Furthermore, Figures 2(b) and 2(d) corresponding to NTK (with Erf and ReLU respectively) do not exhibit significantly different values from the NNGP counterparts. These observations empirically verify our theoretical results.
As $d_{0}$ increases, $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$ increases at similar rates for $Q_{GP-Erf}$ and $\Theta_{NTK-Erf}$. With ReLU activation, $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$ remains almost constant for $Q_{GP-ReLU}$ and exhibits an increasing trend for $\Theta_{NTK-ReLU}$ — implying less collapse for NTK. All these observations corroborate our theory on the limitation of analyzing NC with NTK. We present additional experimental results for imbalanced datasets in Appendix H and show that for a given $N$, the trends of $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$ for increasing $d_{0}$ can vary based on the imbalance ratio of classes (i.e., $n_{1}/n_{2}$). Nevertheless, the trends in $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$ for NNGP with increasing $d_{0}$, resemble that of a trained 2L-FCN with a large hidden layer width $d_{1}=2000$ (see Figure 3(b) corresponding to the Erf activation). These observations provide zero-order reasoning for NN behavior where the feature mapping is learned based on data properties such as dimension.
@@ 5.4 Experiments with High-Dimensional Gaussian Data
Figure 2: $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$ of the post-activation NNGP kernel ($Q^{(1)}_{GP}$) and NTK ($\Theta^{(2)}_{NTK}$) for Erf and ReLU activations on dataset $\mathcal{D}_{1}(N,d_{0})$ (as per equation 17).
6 Activation Variability in the Feature Learning Regime
@@
The explicit kernel formulations in the infinite width limit $(d_{1}\to\infty)$ have allowed us to go beyond the unconstrained features assumption and preserve a link between the features and the data. Yet, the NC phenomenon relates to NN training, while NNGP relates to NN at initialization and NTK has been found unsuitable for NC analysis. Thus, we wish to contrast NNGP with a kernel that is an alternative to NTK and takes into account both optimization and data. To this end, we transition to a large but finite width ($d_{1}\gg 1$) and large sample ($N\gg 1$) setting and analyze the recently introduced ‘adaptive kernels’ approach by Seroussi et al. [40] for fully connected networks.
6.1 Equations of State (EoS)
@@
A transition from the infinite to finite width regime can introduce various corrections to the pre-and post-activations of a $L$-layer FCN. In this context, Seroussi et al. [40] have observed the following dominant corrections: (1) The mean and covariance of the pre-activations deviate from that of a random FCN and, (2) the collective effect of activations from the $(l+1)^{th}$ and $(l-1)^{th}$ layers determine the covariance of activations in the $l^{th}$ layer. Based on these observations, they employ a Variational Gaussian Approximation (VGA) approach to propose the following system of equations for the pre and post-activation kernels $K^{(l)}(\cdot,\cdot),Q^{(l)}(\cdot,\cdot),l\in[L]$ respectively. We formally define the EoS for a 2-layer FCN as follows (based on specializing the generic $L$-layer formulation in equation 5 in [40] to $L=2$, as done in equation 95 in their arxiv extended version):
@@
$\bullet$ Relationship with NNGP: At initialization, we set $\mathbf{C}=(\sigma_{w}^{2}/d_{0})\mathbf{I}$, which implies that $\mathbf{K}^{(1)}=(\sigma_{w}^{2}/d_{0})\mathbf{X}^{\top}\mathbf{X}$. This resulting $\mathbf{K}^{(1)}$ exactly matches the kernel matrix for the pre-activation GP kernel (10) $K_{GP}^{(1)}(\mathbf{x}^{c,i},\mathbf{x}^{c^{\prime},j})=(\sigma_{w}^{2}/d_{0})\mathbf{x}^{c,i\top}\mathbf{x}^{c^{\prime},j}$ as $\sigma_{b}\to 0$. Similarly, the matrix $\mathbf{Q}^{(1)}$ corresponds to the $Q_{GP-Erf}(\cdot,\cdot)$ kernel function defined in (11). The EoS provides a mechanism for transitioning from NNGP kernels to finite-width-based kernels that adapt to the data. Intuitively, observe that the predictions $\overline{\mathbf{f}}$ are formulated based on kernel ridge regression with $\mathbf{Q}^{(1)}$ and $\mathbf{y}$. The $\mathbf{Q}^{(1)}$ matrix along with $\mathbf{y}$ and the initial predictions $\overline{\mathbf{f}}$ are then used to update the weight covariance matrix $\mathbf{C}$. Notice that every entry $[\mathbf{C}^{-1}]_{ij}$ involves a trace operation on a matrix product, resulting in a weighted sum across entries of $\partial_{C_{ij}}\mathbf{Q}^{(1)}$ (i.e the $N^{2}$ pairs of data samples).
$\bullet$ Finite width corrections of EoS on NC1: We solve the EoS (initialized with $\mathbf{C}=(\sigma_{w}^{2}/d_{0})\mathbf{I}_{d_{0}}$) and obtain the stable state using the Newton-Krylov method with an annealing schedule, as originally proposed by [40]. (see Appendix G for details). At initialization, $\mathbf{Q}^{(1)}$ is exactly described by the limiting NNGP kernel matrix, which has been analyzed in the previous section. Now, by solving the EoS with the final annealing factors as 20002000 20002000 and 500500 500500 (which correspond to a 2L-FCN with hidden layer widths $d_{1}=2000,500$ respectively), we illustrate the NC1 metrics of $\mathbf{Q}^{(1)}$ in Figure 3 for the running example of a balanced 2 class dataset $\mathcal{D}_{1}(N,d_{0})$ (17). Notice that for $d_{1}=2000$, the $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$ values for $\mathbf{Q}^{(1)}$ in Figure 3(a) closely resemble the plots for the limiting NNGP kernel $Q_{GP-Erf}$ in Figure 2(a) and the NTK in Figure 2(b). However, we observe noticeable changes in the metrics when $d_{1}=500$. Especially, for $d_{0}\geq 8$ and $N\geq 512$, the $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$ values for the EoS in Figure 3(c) exhibit a noticeable reduction compared to $Q_{GP-Erf}$ in Figure 2(a) and $\Theta_{NTK-Erf}$ in Figure 2(b). Based on this ‘kernel vs. kernel’ analysis, the reduction in $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$ values for EoS reflect the departure of $\mathbf{Q}^{(1)}$ from the initial NNGP state to a feature learning state (based on finite widths).
@@
(b) 2L-FCN $d_{1}=2000$
(c) EoS $d_{1}=500$
(d) 2L-FCN $d_{1}=500$
Figure 3: $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$ of the $\mathbf{Q}^{(1)}$ kernel obtained by solving the EoS (a), (c) and 2L-FCN (b), (d) for the same $d_{1}$ on dataset $\mathcal{D}_{1}(N,d_{0})$ (as per equation 17).
6.2 Activation Variability with Adaptive Kernels (EoS) and 2L-FCN
@@
Setup: We train a 2L-FCN with $d_{1}=500,\sigma_{w}=1,\sigma_{b}=0$ and Erf activation using (vanilla) Gradient Descent with a learning rate of $10^{-3}$ and weight-decay $10^{-6}$ for 10001000 10001000 epochs on datasets described below. For EoS, we employ the same setup described above, with the final annealing factor of $d_{1}=500$ and $\sigma_{a}^{2}=1/128$ (as per the critical scaling value [40]). Similar to the formulation of $\mathcal{D}_{1}(N,d_{0})$ for $C=2$, we formulate $\mathcal{D}_{2}(N,d_{0})$ for $C=4$ as follows:
$\bullet$ $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$ of EoS provides a good approximation of FCN. Let us consider the running example with the balanced 2 class dataset $\mathcal{D}_{1}(N,d_{0})$ for varying $N,d_{0}$ as per (17). The EoS primarily aims to capture the finite-width corrections (as discussed above) depending on the scaling of $N,d_{0}$, and $d_{1}$. To this end, observe from Figure 3(c) that for $d_{1}=500$, the trends of $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$ for EoS vary depending on the scale of $N,d_{0}$. Figure 3(d) illustrates the the actual 2L-FCN behavior. For $d_{0}=\{1,2\}$, although $N$ is scaled to larger values, the $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$ of EoS results in similar values and resemble the 2L-FCN case. However, as the input dimension increases to $d_{0}=\{8,32\}$, the larger value of $N=\{1024\}$ tends to deviate from the 2L-FCN behaviour. Similar deviations were observed even after choosing higher/lower learning rates for 2L-FCN (i.e $5*10^{-3},2*10^{-3},5*10^{-4},10^{-4}$) and weight-decays (ex: $10^{-5},10^{-4}$). Nonetheless, for the higher dimension of $d_{0}=128$ notice that the larger values of $N=\{512,1024\}$ are required for the EoS to match the 2L-FCN. Additionally, by training deeper FCN’s with $L=\{3,4,5,6\}$ layers and hidden layer widths 500500 500500 on $\mathcal{D}_{1}(N,d_{0})$, we observed that the EoS trends (which correspond to 2L-FCN) can also be used to estimate the trends of $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$ reduction in the penultimate layers of these networks (see Figure 14 in Appendix H).
The trends in $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$ for EoS and 2L-FCN hold even for datasets $\mathcal{D}_{2}(N,d_{0})$ with $C>2$. Observe from Figure 4 that the $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$ values for the kernels $Q_{GP-Erf},\Theta_{NTK-Erf}$, EoS and for 2L-FCN are consistently lower than the $C=2$ case as shown in Figures 2(a), 2(b), 3(c), 3(d), even when comparing them at $d_{0}=1$. We believe this is because $\mathcal{D}_{2}(N,d_{0})$ is constructed by adding 22 22 new classes to $\mathcal{D}_{1}(N,d_{0})$ whose distance between the means is much larger than the within class co-variances.
$\bullet$ On the effects of class imbalances: One of the key conditions for the EoS to approximate the $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$ behavior of a 2L-FCN is the presence of sufficient data samples [40] (depending on the scale of the input dimension $d_{0}$ as shown above). Thus, extreme class imbalances can lead to biased $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$ trends in EoS. Consider a collection of imbalanced datasets based on $\mathcal{D}_{1}(N,d_{0})$ where $N=2048$ is split into two classes as follows: Case 1: $(768,1280)$, Case 2: $(512,1536)$, Case 3: $(256,1792)$. We observe that for Case 3 where the imbalance ratio is relatively large, the $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$ values in EoS are slightly larger than 2L-FCN (Figure 9 in Appendix H). On the other hand, the imbalance ratios in Case 1 and Case 2 lead to good approximations with 2L-FCN. A similar observation can be made for the dataset $\mathcal{D}_{2}(N,d_{0})$ with $C=4$ and $N=1024$ from Figure 10.
$\bullet$ Implications: Our analysis showcases the dependence of $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$ on the activation functions and indicates that an increase in the data complexity by increasing the dimension of the data typically leads to larger $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$ values. Furthermore, the $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$ values depend heavily on the relative positions of the data points. In particular, when the means $\{\mu_{c}\}$ are well separated with smaller $\{\sigma_{c}\}$, even a large number of classes can lead to smaller $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$ values. These observations explain the empirical results by Papyan et al. [3] (Figure 6) where extensive experiments on complex datasets (e.g., ImageNet) led to relatively less collapse (i.e larger $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$) than simpler datasets such as MNIST. We present a broader discussion on the limitations of our work and future efforts in Appendix J, and a discussion on formulating NC1 metrics that consider the structure of the data in Appendix F.
@@
Figure 4: $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$ of the limiting kernels, adaptive kernel (EoS) with final annealing factor 500500 500500 and 2L-FCN with $d_{1}=500$ and Erf activation on dataset $\mathcal{D}_{2}(N,d_{0})$ (as per equation 19).

### v3

@@ 5.3 On the Limitations of NTK to Exhibit ‘More’ Collapsed Features than NNGP
Under Assumptions 1-3, let $\phi(\cdot)$ be the ReLU activation. Denote by $\mathbf{H}_{GP},\mathbf{H}_{NTK}$ the features associated with NNGP $Q_{GP-ReLU}^{(1)}$ and NTK $\Theta_{NTK-ReLU}^{(2)}$, respectively. Then:
@@
Appendix D presents the proof by calculating the expected values of $Q_{GP-ReLU}^{(1)}(\mathbf{x}^{c,i},\mathbf{x}^{c^{\prime},j})$ and employing Theorem 4.1. For a better understanding of the result, consider the balanced class scenario with $n_{1}=n_{2}=N/2$. This gives us: $\mathbb{E}\left[\mathcal{N}\mathcal{C}_{1}(\mathbf{H}_{GP/NTK})\right]=2(\sigma_{1}^{2}+\sigma_{2}^{2})/(\mu_{1}-\mu_{2})^{2}+\Delta_{h.o.t}$, which captures the sum of the within-class variance of $\mathbf{X}$ in the numerator and the between-class variance in the denominator of the first term.
The Erf activation case.
@@
We present a similar analysis with the NNGP and NTK with Erf activation in Appendix E (as the terms involved in the formulation are relatively complex than the ReLU case). In the case of ReLU with balanced data, observe that the numerator (corresponding to $\Sigma_{W}(\mathbf{H}_{GP/NTK})$) solely depends on $\sigma_{c}^{2}$, however, for Erf activation, our analysis shows a dependence on terms $\propto\sigma_{c}^{2}\mu_{c}^{-6}$ i.e, it depends on inverse of higher powers of class means $\mu_{c}$ as well (see (92) in Appendix E). Similar analysis for $\boldsymbol{\Sigma}_{B}(\mathbf{H}_{GP/NTK})$ with Erf in (100) shows a dependence on terms $\propto\sigma_{c}^{2}\mu_{c}^{-4}$. Importantly, under Assumptions 1-3, we show similar values of the expected NC1 metric for NNGP and NTK even for the Erf activation.
Key Takeaway.
@@
Our results show that even with 1-D Gaussian data, the expected $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$ of the NTK closely approximates the NNGP counterparts. Perhaps surprisingly, this shows that NTK does not represent more collapsed features than NNGP, despite being associated with NN gradient-based optimization. Namely, we have established another result that shows that training in the lazy regime provably deviates from the practical feature learning of NNs (Woodworth et al., 2020; Ghorbani et al., 2020; Wei et al., 2019; Yehudai & Shamir, 2019; Li et al., 2020).
5.4 Experiments with High Dimensional Data
@@
Figure 2: $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$ of (a) the post-activation NNGP kernel ($Q^{(1)}_{GP-Erf}$), (b) NTK ($\Theta^{(2)}_{NTK-Erf}$), (c) 2L-FCN with $d_{1}=500$ for Erf activation on dataset $\mathcal{D}_{1}(N,d_{0})$ (as per (17)).
(a) $Q_{GP-ReLU}$
(b) $\Theta_{NTK-ReLU}$
(c) 2L-FCN $d_{1}=500$
Figure 3: $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$ of (a) the post-activation NNGP kernel ($Q^{(1)}_{GP-ReLU}$), (b) NTK ($\Theta^{(2)}_{NTK-ReLU}$), (c) 2L-FCN with $d_{1}=500$ for ReLU activation on dataset $\mathcal{D}_{1}(N,d_{0})$ (as per (17)).
Setup.
We conduct experiments on datasets with varying sample sizes and input dimensions to verify our theoretical results and show that insights generalize (e.g., beyond $d_{0}=1$). For $C=2$, a dataset size $N$ chosen from $\{128,256,512,1024\}$, and input dimension $d_{0}$ chosen from $\{1,2,8,32,128\}$, we create the data vector and label pairs as follows:
The vectors and labels from the dataset can then be arranged into the matrix form (as described in the setup) for analysis. The sampling procedure is repeated 1010 1010 times for each $(N,d_{0})$.
Erf activation leads to more collapsed features than ReLU.
For the low-dimensional case of $d_{0}=1$ and Erf activation, observe from Figure 2(a) that $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$ has a small value of $\approx 10^{-2.1}$. On the contrary, Figure 3(a) illustrates that for ReLU, $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$ is more than an order of magnitude larger ($\approx 10^{-0.95}$) than the former. Furthermore, Figures 2(b) and 3(b) corresponding to NTK (with Erf and ReLU respectively) do not exhibit significantly different values from the NNGP counterparts. These observations empirically verify our theoretical results above for $d=1$. Furthermore, notice that the $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$ values are higher in the case of ReLU than Erf across all the dimensions. We also observe such behavior when training the 2L-FCN networks with hidden layer width $d_{1}=500$ (see Appendix G for training details). When $d_{1}=500$, the finite width effects of training come into play (see following sections) and we observe a shift in the trend of $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$ for $d_{0}>32$. Nonetheless, the Erf case in Figure 2(c) still exhibits more collapse than ReLU case in Figure 3(c).
NTK does not result in more collapsed features than NNGP in higher dimensions.
As the data dimension $d_{0}$ increases, $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$ increases at similar rates for $Q_{GP-Erf}$ and $\Theta_{NTK-Erf}$. With ReLU activation, $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$ remains almost constant for $Q_{GP-ReLU}$ and exhibits an increasing trend for $\Theta_{NTK-ReLU}$implying less collapse for NTK. All these observations corroborate our theory on the limitation of analyzing NC with NTK. We present additional experimental results for imbalanced datasets in Appendix G and show that for a given $N$, the trends of $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$ for increasing $d_{0}$ can vary based on the imbalance ratio of classes (i.e., $n_{1}/n_{2}$). These observations provide zero-order reasoning for NN behavior where the feature mapping is learned based on data properties such as dimension.
6 Activation Variability of data-aware GP Kernels
@@
The explicit kernel formulations in the infinite width limit $(d_{1}\to\infty)$ have allowed us to go beyond the unconstrained features assumption and preserve a link between the features and the data. Yet, the NC phenomenon relates to NN training, while NNGP relates to NN at initialization and NTK has been found unsuitable for NC analysis. Thus, we wish to contrast NNGP with a kernel that is an alternative to NTK and takes into account both optimization and data. To this end, we transition to a large but finite width ($d_{1}\gg 1$) and large sample ($N\gg 1$) setting and analyze the data-aware GP kernels by (Seroussi et al., 2023) for FCNs.
6.1 Equations of State (EoS) for the data-aware GP Kernel
@@
A transition from the infinite to finite width regime can introduce various “corrections” to the pre-and post-activations of a $L$-layer FCN, which account for the different behavior. In this context, (Seroussi et al., 2023) have observed the following dominant corrections: (1) The mean and covariance of the pre-activations deviate from that of a random FCN in a nearly Gaussian manner and, (2) the collective effect of activations from the $(l+1)^{th}$ and $(l-1)^{th}$ layers determine the covariance of activations in the $l^{th}$ layer. Therefore, while for infinite-width DNNs, the kernels representations are inert, for finite DNNs they adapt to the data and yield a tractable data-aware Gaussian Process. Specifically, based on the aforementioned observations, (Seroussi et al., 2023) employ a Variational Gaussian Approximation (VGA) approach and derived the following system of equations for the pre and post-activation kernels $K^{(l)}(\cdot,\cdot),Q^{(l)}(\cdot,\cdot),l\in[L]$ respectively. The solution of this system models the state of the DNN at convergence (after the training phase), and has been empirically shown to yield accurate predictions in various settings. We formally define the EoS for a 2-layer FCN as follows (based on specializing the generic $L$-layer formulation in equation 5 in (Seroussi et al., 2023) to $L=2$, as done in equation 95 in their arxiv extended version):
@@
$\bullet$ Relationship with NNGP.
At initialization, we set $\mathbf{C}=(\sigma_{w}^{2}/d_{0})\mathbf{I}$, which implies $\mathbf{K}^{(1)}=(\sigma_{w}^{2}/d_{0})\mathbf{X}^{\top}\mathbf{X}$. This resulting $\mathbf{K}^{(1)}$ exactly matches the kernel matrix for the pre-activation GP kernel (10) $K_{GP}^{(1)}(\mathbf{x}^{c,i},\mathbf{x}^{c^{\prime},j})=(\sigma_{w}^{2}/d_{0})\mathbf{x}^{c,i\top}\mathbf{x}^{c^{\prime},j}$ as $\sigma_{b}\to 0$. Similarly, the matrix $\mathbf{Q}^{(1)}$ corresponds to the $Q_{GP-Erf}(\cdot,\cdot)$ kernel function defined in (11). The EoS provides a mechanism for transitioning from NNGP kernels to finite-width-based kernels that adapt to the data. Intuitively, observe that the predictions $\overline{\mathbf{f}}$ are formulated based on kernel ridge regression with $\mathbf{Q}^{(1)}$ and $\mathbf{y}$. The $\mathbf{Q}^{(1)}$ matrix along with $\mathbf{y}$ and the initial predictions $\overline{\mathbf{f}}$ are then used to update the weight covariance matrix $\mathbf{C}$. Here every entry $[\mathbf{C}^{-1}]_{ij}$ involves a trace operation on a matrix product, resulting in a weighted sum across entries of $\partial_{C_{ij}}\mathbf{Q}^{(1)}$ (i.e the $N^{2}$ pairs of data samples).
6.2 Numerical Solutions of EoS
@@
(b) EoS $d_{1}=500$
Figure 4: $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$ of the $\mathbf{Q}^{(1)}$ kernel obtained by solving the EoS (a) $d_{1}=2000$ (b) $d_{1}=500$ on $\mathcal{D}_{1}(N,d_{0})$.
We solve the EoS using the Newton-Krylov method with an annealing schedule (as originally proposed by (Seroussi et al., 2023)) using the scipy.optimize.newton_krylov python API. We initialize $\mathbf{C}$ with the GP limit value of $(\sigma_{w}^{2}/d_{0})\mathbf{I}_{d_{0}}$ and choose a large annealing factor (ex: $10^{5}$) as the value for $d_{1}$. The result of optimizing with newton_krylov is a new $\mathbf{C}$, which in addition to a lower annealing factor is used as an input for the next newton_krylov function call. This loop is repeated until the end of an annealing schedule. For instance, to analyze the EoS corresponding to $d_{1}=500$, we choose the following step-wise annealing factors:
6.3 Experiments
@@
Setup: We train a 2L-FCN with $d_{1}=500,\sigma_{w}=1,\sigma_{b}=0$ and Erf activation using (vanilla) Gradient Descent with a learning rate of $10^{-3}$ and weight-decay $10^{-6}$ for 10001000 10001000 steps. To numerically solve the EoS, we employ the approach described in Section 6.2, with the final annealing factor of $d_{1}=500$ and $\sigma_{a}^{2}=1/128$ (as per the critical scaling value suggested in (Seroussi et al., 2023)).
NC1 of EoS on linearly separable datasets.
We solve the EoS (initialized with $\mathbf{C}=(\sigma_{w}^{2}/d_{0})\mathbf{I}_{d_{0}}$) and obtain the stable state using the Newton-Krylov method with an annealing schedule, as described above. At initialization, $\mathbf{Q}^{(1)}$ is exactly described by the limiting NNGP kernel matrix, which has been analyzed in the previous section. Now, by solving the EoS with the final annealing factors as 20002000 20002000 and 500500 500500 (which correspond to a 2L-FCN with hidden layer widths $d_{1}=2000,500$ respectively), we illustrate the NC1 metrics of $\mathbf{Q}^{(1)}$ in Figure 4 for the running example of parameterized balanced 2 class datasets $\mathcal{D}_{1}(N,d_{0})$ (17). Notice that for $d_{1}=2000$, the $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$ values for $\mathbf{Q}^{(1)}$ in Figure 4(a) closely resemble the plots for the limiting NNGP kernel $Q_{GP-Erf}$ in Figure 2(a) and the NTK in Figure 2(b). However, we observe noticeable changes in the metrics when $d_{1}=500$. Especially, for $d_{0}\geq 8$ and $N\geq 512$, the $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$ values for the EoS in Figure 4(b) exhibit a noticeable reduction compared to $Q_{GP-Erf}$ and $\Theta_{NTK-Erf}$. Furthermore, notice that such a reduction aligns with the behavior of 2L-FCN in Figure 2(c) and reflects the departure of $\mathbf{Q}^{(1)}$ in EoS from the initial NNGP state to a feature learning state.
NC1 of EoS on non-linearly separable datasets.
Unlike the separable Gaussian datasets in the above sections, the real-world datasets are relatively more complex. To simulate such a scenario, we sample $\mathbf{x}^{1,i}\sim\mathcal{N}(-2*\mathbf{1}_{d_{0}},4*\mathbf{I}_{d_{0}}),y^{1,i}=-1,i\in[N/2]$ for class 11 11 and $\mathbf{x}^{2,j}\sim\mathcal{N}(2*\mathbf{1}_{d_{0}},4*\mathbf{I}_{d_{0}}),y^{2,j}=1,j\in[N/2]$ for class 22 22 of our dataset. Essentially, these are scenarios where there is a significant overlap between samples of the two classes. First, we note that we had to increase the learning rate of our 2L-FCN from $10^{-3}$ to $5\cdot 10^{-3}$ and run GD for 20002000 20002000 epochs for convergence. For dimensions $d_{0}=\{8,16,32\}$, the EoS reasonably approximates the 2L-FCN but for $d_{0}=\{64,128\}$, the $\mathcal{N}\mathcal{C}_{1}(\mathbf{H})$ values for 2L-FCN turned out to be almost twice as large as the EoS (see Figure 5). We also present additional experiments with more number of classes, imbalanced classes and deeper FCN networks in Appendix G.
@@
7 Limitations
We highlight the difficulties in the theoretical/empirical analysis of NC1 with EoS. The primary bottleneck is a lack of rigorous study on the existence and uniqueness of solutions (As also highlighted by (Seroussi et al., 2023)). Since we deviate from the lazy regime and deal with kernels in the feature learning setup, we cannot expect simpler closed-form solutions like the limiting NNGP/NTK for the EoS. However, numerical solutions to the EoS can sometimes be time-consuming and require a manual selection of the annealing schedule. This is a tradeoff that can be improved with future research. Furthermore, the role of scaling $N,d_{0},d_{1}$ on NC1 is yet to be fully understood and we hope that our analysis lays the groundwork for such efforts.
Remark.